OpenStudy (anonymous):

explain how to factor 4x^2-4x-48, explain how your factors could be check accuracy?

OpenStudy (anonymous):

4(x-4)(x+3)

OpenStudy (dumbcow):

step 1: factor out a 4 4(x^2 - x - 12) step 2: find factors of neg. 12 that add up to neg. 1 -4*3 = -12 -4+3 = -1 step 3: put in factor form 4(x-4)(x+3)